OpenStudy (anonymous):

solve for h if 5gh=f ????

OpenStudy (sandra):

hmmm, I think it's just asking you to isolate h on one side of the equation?

OpenStudy (sandra):

so 1. 5gh/5 = f/5 2. gh/g = f/5 /g 3. h = f/5g

OpenStudy (sandra):

does that sound right?

OpenStudy (anonymous):

yes thank you